# RoboUniView: Visual-Language Model with Unified View Representation for Robotic Manipulaiton

### v1

@@ Abstract
Utilizing Vision-Language Models (VLMs) for robotic manipulation represents a novel paradigm, aiming to enhance the model’s ability to generalize to new objects and instructions. However, due to variations in camera specifications and mounting positions, existing methods exhibit significant performance disparities across different robotic platforms. To address this challenge, we propose RoboUniView in this paper, an innovative approach that decouples visual feature extraction from action learning. We first learn a unified view representation from multi-perspective views by pre-training on readily accessible data, and then derive actions from this unified view representation to control robotic manipulation. This unified view representation more accurately mirrors the physical world and is not constrained by the robotic platform’s camera parameters. Thanks to this methodology, we achieve state-of-the-art performance on the demanding CALVIN benchmark, enhancing the success rate in the $D\to D$ setting from 88.7% to 96.2%, and in the $ABC\to D$ setting from 82.4% to 94.2%. Moreover, our model exhibits outstanding adaptability and flexibility: it maintains high performance under unseen camera parameters, can utilize multiple datasets with varying camera parameters, and is capable of joint cross-task learning across datasets. Code is provided for re-implementation. https://github.com/liufanfanlff/RoboUniview
@@ 4.2 Comparisons with Other State-of-the-Art Methods
Imitation Performance. We fine-tune RoboUniView using the demonstrations from the Split D training set, and evaluate its imitation performance on episodes sampled from Split D ($D\to D$). It takes about 2 days for training on 8 NVIDIA 80G A100 GPUs. As shown in Table 1, RoboUniView significantly outperforms all methods across all metrics. The success rate of task1 is improved from 0.887 to 0.962. Even more impressive, in sequence of consecutive tasks, RoboUniView increase the success rate of task5 from 0.349 to 0.563 and raise the average successful sequence length from 2.968 to 3.855. This result is particularly commendable as the complexity and challenge of subsequent tasks significantly increase with the progression of the tasks. This primarily stems from the fact that the initial state of each subsequent task is heavily dependent on the completion state of the previous task, leading to increasingly diverse starting conditions.
Zero-Shot Generalization. We also fine-tune RoboUniView on the ABC split and test on the D split ($ABC\to D$), where the D split presents a completely different visual environment from ABC. It takes about 5 days for training on 8 NVIDIA 80G A100 GPUs. As shown in Table 1, RoboUniView improves the success rate of task1 from 0.824 to 0.942, and the average successful sequence length from 2.47 to 3.647, compare to best method. It demonstrate RoboUniView’s strong capability in zero-shot generalization.

### v2

@@ Abstract
Utilizing Vision-Language Models (VLMs) for robotic manipulation represents a novel paradigm, aiming to enhance the model’s ability to generalize to new objects and instructions. However, due to variations in camera specifications and mounting positions, existing methods exhibit significant performance disparities across different robotic platforms. To address this challenge, we propose RoboUniView in this paper, an innovative approach that decouples visual feature extraction from action learning. We first learn a unified view representation from multi-perspective views by pre-training on readily accessible data, and then derive actions from this unified view representation to control robotic manipulation. This unified view representation more accurately mirrors the physical world and is not constrained by the robotic platform’s camera parameters. Thanks to this methodology, we achieve state-of-the-art performance on the demanding CALVIN benchmark, enhancing the success rate in the $D\to D$ setting from 93.0% to 96.2%, and in the $ABC\to D$ setting from 92.2% to 94.2%. Moreover, our model exhibits outstanding adaptability and flexibility: it maintains high performance under unseen camera parameters, can utilize multiple datasets with varying camera parameters, and is capable of joint cross-task learning across datasets. Code is provided for re-implementation. https://github.com/liufanfanlff/RoboUniview
@@ 4.2 Comparisons with Other State-of-the-Art Methods
Imitation Performance. We fine-tune RoboUniView using the demonstrations from the Split D training set, and evaluate its imitation performance on episodes sampled from Split D ($D\to D$). It takes about 2 days for training on 8 NVIDIA 80G A100 GPUs. As shown in Table 1, RoboUniView significantly outperforms all methods across all metrics. The success rate of task1 is improved from 0.930 to 0.962. Even more impressive, in sequence of consecutive tasks, RoboUniView increase the success rate of task5 from 0.400 to 0.563 and raise the average successful sequence length from 3.300 to 3.855. This result is particularly commendable as the complexity and challenge of subsequent tasks significantly increase with the progression of the tasks. This primarily stems from the fact that the initial state of each subsequent task is heavily dependent on the completion state of the previous task, leading to increasingly diverse starting conditions.
Zero-Shot Generalization. We also fine-tune RoboUniView on the ABC split and test on the D split ($ABC\to D$), where the D split presents a completely different visual environment from ABC. It takes about 5 days for training on 8 NVIDIA 80G A100 GPUs. As shown in Table 1, RoboUniView improves the success rate of task1 from 0.922 to 0.942, and the average successful sequence length from 3.270 to 3.647, compare to best method. It demonstrate RoboUniView’s strong capability in zero-shot generalization.